OpenStudy (aravindg):

A homogeneous chain of length l lies on a table. What is the maximum length l, of the chain that can hang over the table if the coefficient of friction is µ?

OpenStudy (anonymous):

\[\mu l/(\mu +1 )\] i think so

OpenStudy (anonymous):

The above is correct. The length is a maximum when the force due to gravity on the hanging chain just overcomes the force due to friction of the chain on the table. F = CF . m(table) . g = m(hanging). g mass is proportional to length do substitute for L CF. L(t).g = L(h).g And L = L(t) + L(h) So CF (L - L(h)).g = L(h).g CF.L -CF.L(h) = L(h) L(h) = CF.L/ (1+ CF)